# KAN we improve on HEP classification tasks? Kolmogorov-Arnold Networks applied to an LHC physics example

### v1

@@ Abstract
Recently, Kolmogorov-Arnold Networks (KANs) have been proposed as an alternative to multilayer perceptrons, suggesting advantages in performance and interpretability. We study a typical binary event classification task in high-energy physics including high-level features and comment on the performance and interpretability of KANs in this context. We find that the learned activation functions of a one-layer KAN resemble the log-likelihood ratio of the input features. In deeper KANs, the activations in the first KAN layer differ from those in the one-layer KAN, which indicates that the deeper KANs learn more complex representations of the data. We study KANs with different depths and widths and we compare them to multilayer perceptrons in terms of performance and number of trainable parameters. For the chosen classification task, we do not find that KANs are more parameter efficient. However, small KANs may offer advantages in terms of interpretability that come at the cost of only a moderate loss in performance.
@@ 1 Introduction
Classifying events as signal or background is a crucial ingredient of data analysis at collider experiments. At the Large Hadron Collider (LHC), separating small signals from large backgrounds is an omnipresent challenge. To achieve higher precision in the analysis of collider data, excellent classifiers are necessary. Machine-learning-based classifiers have a long history in high-energy physics (HEP). For example, the observation of electroweak production of single top quarks in 2009 at the Tevatron [1, 2] was aided by boosted decision trees and by shallow neural networks, i.e. multilayer perceptrons (MLPs) with one hidden layer. With the development of deep neural networks, MLPs with several hidden layers have been proposed for HEP classification tasks [3] and have become a standard tool for event classification, particle identification, fast simulations and many more applications at the LHC [4, 5, 6, 7, 8].
@@
The strong performance of MLPs comes with a trade-off in terms of interpretability. This is in particular true for deep MLPs with their large number of trainable parameters. While many methods exist that help to interpret the output of MLPs for given input examples [9], interpretability, i.e. the “ability to explain or to present in understandable terms to a human” [10], is an intrinsic property of the model and remains a challenge for MLPs. However, understanding what such a model has learned about the underlying physics should be of genuine interest in physics applications.
@@
Recently, Kolmogorov-Arnold Networks (KANs) have been proposed as an alternative to MLPs [11]. While MLPs are connected to the universal approximation theorem [12], KANs are motivated by the Kolmogorov-Arnold representation theorem [13]. Practically, KAN layers have learnable activation functions on the edges that are summed on the nodes. In contrast, MLP layers have learnable weights on the edges that are used as the input to fixed activation functions on the nodes. While networks based on the Kolmogorov-Arnold representation theorem were proposed before [14, 15, 16, 17, 18, 19, 20, 21], recently the capability of KANs in terms of performance and interpretability was highlighted [11]. In Ref. [11], KANs were found to have promising performance with a substantially smaller number of trainable parameters than MLPs. In addition, the potential for interpretability by approximating the learned activation functions with a set of known functions was discussed. Ref. [11] has sparked active discussion on the potential advantages of KANs and their relation to MLPs [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73].
@@
We apply KANs to a typical HEP event classification task. As an example, we choose the binary separation of the associated production of a Higgs boson with a single top quark ($tH$) and with a top quark and an anti-top quark ($t\bar{t}H$) at the LHC, where the Higgs boson decays to a pair of photons ($H\to\gamma\gamma$). We study the interpretability of KANs for this classification task. Additionally, we compare KANs to MLPs in terms of performance and parameter efficiency, where we use KANs and MLPs with different numbers of layers and nodes per layer. In addition, we document our findings in the practical training of KANs. To our knowledge, this is the first application of KANs to a task in particle physics.
@@ 2 Kolmogorov-Arnold Networks
For the comparison to KANs, we briefly summarize the concept of MLPs. An MLP consists of multiple layers of nodes, each connected to nodes in subsequent layers through weighted edges. The core component of an MLP is the fully connected layer, which holds the trainable parameters defining the strength of the connections between nodes of two layers. Each layer applies a linear transformation, represented by a weight matrix $\mathbf{W}$ and a bias vector $\vec{b}$, followed by an activation function $\mathcal{A}$. The transformation applied in each MLP layer can then be written as $\vec{y}=\mathcal{A}\left(\mathbf{W}\vec{x}+\vec{b}\right)$, where $\vec{x}$ denotes the input to the layer and $\vec{y}$ is its output. The activation function introduces non-linearity in the model and is a hyperparameter that has to be chosen. Common choices include the rectified linear unit $\mathrm{ReLU}(x)=\max{(0,x)}$, the logistic sigmoid function $\sigma(x)$, and the hyperbolic tangent function.
In contrast, KANs are inspired by the Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as a finite sum of continuous functions of only one variable. Formally, for any continuous real-valued function $f(x_{1},x_{2},\ldots,x_{n})$, continuous functions $\phi_{i(j)}$ exist such that
@@
Motivated by the theorem, the appropriate network architecture to approximate a multivariate function of $n$ variables consists of two layers with $n$ input nodes, $12n+$ 1 hidden nodes, and a single output node. However, the authors of Ref. [11] generalized the concept by defining a KAN-layer as a basic building block. As in MLPs, the number of nodes in these layers can be customized and they can be stacked arbitrarily to enhance the performance of the model. Similar to MLPs, each node in a given layer is connected to each node of the subsequent layer. For each edge, an individual, learnable activation function is used. On the nodes, only the sum operation over all incoming edges is performed.
@@
The weights $w_{1,2}$ and the basis-function coefficients $c_{i}$ are the trainable parameters of the spline. The basis functions $B_{i}$ are chosen as polynomials of degree $k$, with default value $k=3$. The grid parameter $G$ determines how many basis functions build the B-spline and serves as a hyperparameter of the KANs. Furthermore, the domain of the activation function needs to be chosen, which can be updated several times during network training to match the input range of the activation function. Specifically, for given parameters $k$, $G$ and the domain $[t_{0},t_{G}]$, a vector $\vec{t}=(t_{-k},\ldots,t_{0},\ldots,t_{G+k})$ of equidistant knot points is constructed. Then, $G+k$ basis functions $B_{i}^{k}(x)$ are recursively defined:
@@ 4 Results
The output layers of all trained models consist of a single node. Although KANs learn activation functions and a learnable function can also be placed on the output node, we choose the sigmoid function to normalize the model outputs to the range $(0,1)$. Also for the MLPs, we choose the sigmoid function as output activation. For all trainings, we use the binary cross-entropy as loss function. We consider $tH$ events as signal (label 1) and $t\bar{t}H$ as background (label 0).
We use the Adam [90] optimizer to train our models. For the KAN trainings, we also compare with the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) [91] optimizer, as it was used in Ref. [11]. For Adam, we find stable trainings for all tested learning rates in the range $[10^{-4},10^{-3}]$ and select $3\times 10^{-4}$ for all trainings. The training is performed in mini-batches of batch size $256$. Trainings with the L-BFGS optimizer are performed using the entire dataset in full-batch training with a learning rate of $10^{-3}$. No significant performance differences between KANs trained with the two optimizers were found and hence we use Adam for all trainings discussed below due its faster convergence. To ensure that all models converge during training and to allow for a fair performance comparison, we employ early stopping. The loss obtained from the validation dataset is monitored during training and if there is no improvement over 25 epochs, the training is terminated. The model parameters from the epoch with the lowest loss on the validation set are used for the comparisons.
@@
The evolution of the loss and accuracy [^4] The accuracy is defined as the fraction of correctly classified examples with a decision threshold of 0.5 in the network output. over the training is shown in Fig. 3 for three KANs. As expected, the single-layer KAN shows the lowest performance among these models but still reaches an accuracy of $81.5\%$ on the validation set, which is only about $1.5$ percentage points lower than the accuracy of the 2-layer and 4-layer KANs. The latter two models reach similar accuracies and loss values. The two-layer KAN, with its wide second layer, has the highest number of trainable parameters in this comparison. It converges within the fewest number of training epochs and at the same time shows the largest generalization gap. While the range of the input variables is fixed with the min-max scaling, this is not the case for the input range of subsequent layers in the networks. Therefore, we use the default setting of updating the domains ten times within the first 50 epochs of training. Training instabilities are visible at these epochs.
@@
The output distributions of the test dataset classified by these KANs are shown in Fig. 4. The separation of the two classes is clearly visible with the networks accumulating the majority of the events close to the respective label. Also here, only slight differences are visible between the two- and the four-layer network, while the 1-layer KAN achieves a visibly worse separation.
A possible advantage of KANs over MLPs lies in their potential for interpretability. While the patterns learned by MLPs are usually embedded in large matrices of trainable parameters and thus hard to understand, multiple trained KAN parameters can be visualized in form of a single spline. As demonstrated in Ref. [11], patterns learned by KANs can often be understood more easily. However, these examples are of much lower dimension and complexity than typical HEP machine-learning tasks. In our study with 22 input features, we find that the interpretability of wide models, such as the 22–45–1 KAN, is limited. For instance, this particular model consists of more than $1000$ learnable activation functions. Its visualization is hence complex and difficult for humans to interpret. Therefore, we focus on shallow KAN structures for the interpretability.
In Fig. 5, the 1-layer KAN is depicted together with its learned activation functions. The corresponding figure for the KAN with a shallow second layer (22–3–1) is shown in the Appendix. In the following, we discuss the interpretability of the learned activation functions.
@@
These values directly indicate the importance of the different input features for the KAN classification output due to the simple summation of spline output values on the nodes. While also for MLPs, methods can be applied to estimate the importance of certain input features for classification outputs [9], this information is straightforward to obtain for the KAN.
@@
The KAN 22–1 uses a single sum operation over univariate functions of the input features. Therefore, this network is expected to exploit the features with strong discrimination between the two classes. Values of the $L_{1}$-norms close to zero are found for the azimuthal angles, which provide no discrimination power. The largest values of the $L_{1}$-norms are found for the lepton multiplicity, with which the network can identify di-leptonic signatures present in a fraction of the $t\bar{t}H$ events but mostly absent in the $tH$ process, and for variables based on (pseudo-)rapidity differences as well as the jet multiplicity. These reflect the good separation that can be achieved with these variables on their own.
@@
In Fig. 6, the learned activation functions are shown for five examples features. These features were chosen to represent a set of variables with different shapes of their distributions, where the corresponding splines of the single-layer KAN have high $L_{1}$-norms. The distributions of the pre-processed features are also shown for the two classes together with the log-likelihood ratio of the signal over the background. In general, the activation functions learned by the 1-layer KAN are similar to the log-likelihood ratio. While the shape of the activation function often closely follows this ratio, the splines can have a different normalization, as for example visible for the $\eta(j)$ variable. Hence, the 1-layer KAN appears to mainly build a weighted sum of the likelihood ratios of the input variables to achieve the discrimination. The corresponding splines obtained from the training of the 22–3–1 KAN are also shown in the same panels. These differ clearly from the likelihood ratios and the 22–1 KAN splines. This is consistent with the expectation that multi-layer KANs learn more abstract representations of the data. As the splines acting on the same input feature are allowed to have different functional shapes or $L_{1}$-norms, each of the three nodes of the internal layer of the 22–3–1 KAN captures different aspects of the data.
@@
Receiver operating characteristic (ROC) curves for selected models are shown in Fig. 8, including models with deep and shallow structures for both, MLPs and KANs. The overall shape of the ROC curves of KANs and MLPs is very similar, except for the 1-layer KAN, where the area under the curve (AUC) is considerably lower. We observe that an MLP with only a single hidden layer reaches an AUC [^5] The uncertainty in the quoted AUCs from the limited size of the test dataset are approximately $6\times 10^{-4}$. of $0.906$, only slightly below the AUCs obtained by our best MLP ($0.908$) and the best KAN, where the 22–45–1 network also reaches an AUC of $0.908$. We find only slight differences in classification performance between well-tuned KANs and MLPs.
@@
To evaluate the parameter efficiency, we compare the performance of KANs and MLPs as a function of the number of trainable model parameters. A small number of parameters is favorable for a given performance, as smaller models are computationally more efficient, better interpretable (if at all possible), and less prone to overfitting. In Fig. 8, two metrics often used in HEP for classification performance evaluation are included: the AUC, and the background rejection for a fixed signal efficiency, here chosen as $70\,\%$. The rejection is defined as the inverse of the efficiency for a given threshold on the network output. Overall, we find that for very low numbers of parameters, the MLPs outperform the KANs, while for medium and high numbers of parameters, the performance of KANs and MLPs is similar. The smallest MLP trained with only eight nodes in the hidden layer has only $193$ parameters and reaches a background rejection of $12.9$ already [^6] The uncertainty in the quoted rejection values from the limited size of the test dataset is approximately $0.2$.. A similar value is achieved by the 22–3–1 KAN, which has $556$ trainable parameters. While increasing the number of parameters of the MLPs, their rejection saturates at around $14.9$. The KANs achieve rejections ranging from $11.4$ for the single-layer network to $15.2$ for the 22–45–1 KAN. The deeper KANs from this study show a similar performance. Instead of varying the node count of KANs, the number of model parameters can be increased as well by raising the grid parameter $G$. For illustration, the 22–45–1 KAN is included in Fig. 8 when trained with $G=10$ and $G=25$. For KANs with grid parameters considerably higher than the default of $G=5$, we find that these models overtrain faster and generalize worse.

### v2

@@ Abstract
Recently, Kolmogorov-Arnold Networks (KANs) have been proposed as an alternative to multilayer perceptrons, suggesting advantages in performance and interpretability. We study a typical binary event classification task in high-energy physics including high-level features and comment on the performance and interpretability of KANs in this context. Consistent with expectations, we find that the learned activation functions of a one-layer KAN resemble the univariate log-likelihood ratios of the respective input features. In deeper KANs, the activations in the first layer differ from those in the one-layer KAN, which indicates that the deeper KANs learn more complex representations of the data, a pattern commonly observed in other deep-learning architectures. We study KANs with different depths and widths and we compare them to multilayer perceptrons in terms of performance and number of trainable parameters. For the chosen classification task, we do not find that KANs are more parameter efficient. However, small KANs may offer advantages in terms of interpretability that come at the cost of only a moderate loss in performance.
@@ 1 Introduction
Classifying events as signal or background is a crucial ingredient of data analysis at collider experiments. At the Large Hadron Collider (LHC), separating small signals from large backgrounds is an omnipresent challenge. To achieve higher precision in the analysis of collider data, excellent classifiers are necessary. Machine-learning-based classifiers have a long history in high-energy physics (HEP). For example, the observation of electroweak production of single top quarks in 2009 at the Tevatron [1, 2] was aided by boosted decision trees and by shallow neural networks, i.e., multilayer perceptrons (MLPs) with one hidden layer. With the development of deep neural networks, MLPs with several hidden layers have been proposed for HEP classification tasks [3] and have become a standard tool for event classification, particle identification, fast simulations and many more applications at the LHC [4, 5, 6, 7, 8].
@@
The strong performance of MLPs comes with a trade-off in terms of interpretability. Interpretability, i.e., the “ability to explain or to present in understandable terms to a human” [9], remains a challenge for MLPs, particularly for deep networks with many trainable parameters. At the same time, understanding what such a model has learned about the underlying physics is of genuine interest in physics applications. Several methods have been developed to address this challenge of explaining the outputs of MLPs for given input examples [10, 11, 12]. Techniques such as Shapley values [13] and permutation feature importance [14] are established methods for assessing the contribution of individual input features to the model output. Surrogate models, such as LIME [15], aim to explain the reasoning of complex architectures by approximating them with simplified models. On the other hand, approaches like Neural Additive Models [16] provide interpretability by constructing models that are transparent by design.
Recently, Kolmogorov-Arnold Networks (KANs) have been proposed as an alternative to MLPs [17]. While MLPs are grounded in the universal approximation theorem [18], KANs are motivated by the Kolmogorov-Arnold representation theorem [19]. The layers of the KAN have learnable activation functions on the edges that are summed on the nodes. In contrast, MLP layers use learnable weights on the edges as inputs to fixed activation functions on the nodes. Many approaches have been explored to improve the expressiveness and performance of MLPs by introducing learnable activation functions. Examples include implementations based on splines [20, 21], other parametric functions [22, 23, 24, 25] or even neural networks [26] as activation functions.
@@
While networks based on the Kolmogorov-Arnold representation theorem were proposed before [27, 28, 29, 30, 31, 32, 33, 34], recently the capabilities of KANs in terms of performance and interpretability were highlighted [17]. In Ref. [17], KANs were found to have promising performance with a substantially smaller number of trainable parameters than MLPs. KANs offer advantages in terms of interpretability, complementarily to existing explainability methods applicable to both KANs and MLPs, due to their shallower structures with significantly fewer nodes than typical MLPs. Each edge in a KAN contains multiple trainable parameters that determine the shape of a single function. Therefore, an entire KAN can be represented as a comprehensive graph. In addition, the potential for interpretability by approximating the learned activation functions symbolically with a set of known functions was discussed. Ref. [17] has sparked active discussion on the potential advantages of KANs and their relation to MLPs [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86].
@@
We apply KANs to a typical HEP event classification task. As an example, we choose the binary separation of the associated production of a Higgs boson with a single top quark ($tH$) and with a top quark and an anti-top quark ($t\bar{t}H$) at the LHC, where the Higgs boson decays to a pair of photons ($H\to\gamma\gamma$). We study the interpretability of KANs for this classification task. In addition, we compare KANs to MLPs in terms of performance and parameter efficiency, where we use KANs and MLPs with different numbers of layers and nodes per layer. We document our findings in the practical training of KANs. To our knowledge, this is the first application of KANs to a task in particle physics.
@@ 2 Kolmogorov-Arnold Networks
For the comparison to KANs, we briefly summarize the concept of MLPs. An MLP consists of multiple layers of nodes, each connected to nodes in subsequent layers through weighted edges. The core component of an MLP is the fully connected layer, which holds the trainable parameters defining the strength of the connections between nodes of two layers. Each layer applies an affine transformation, represented by a weight matrix $\mathbf{W}$ and a bias vector $\vec{b}$, followed by an activation function $\mathcal{A}$. The transformation applied in each MLP layer can then be written as $\vec{y}=\mathcal{A}\left(\mathbf{W}\vec{x}+\vec{b}\right)$, where $\vec{x}$ denotes the input to the layer and $\vec{y}$ is its output. The activation function introduces non-linearity in the model and is a hyperparameter that has to be chosen. Common choices include the rectified linear unit $\mathrm{ReLU}(x)=\max{(0,x)}$, the logistic sigmoid function $\sigma(x)$, and the hyperbolic tangent function.
In contrast, KANs are inspired by the Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as a finite sum of continuous functions of only one variable. Formally, for any continuous real-valued function $f(x_{1},x_{2},\ldots,x_{n})$, continuous functions $\phi_{i(j)}$ exist, such that
@@
Motivated by the theorem, the appropriate network architecture to approximate a multivariate function of $n$ variables consists of two layers with $n$ input nodes, $12n+$ 1 hidden nodes, and a single output node. However, the authors of Ref. [17] generalized the concept by defining a KAN layer as a basic building block. As in MLPs, the number of nodes in these layers can be customized and layers can be stacked arbitrarily to enhance the performance of the model. Similar to MLPs, each node in a given layer is connected to each node of the subsequent layer. For each edge, an individual, learnable activation function is used. On the nodes, only the sum operation over all incoming edges is performed.
@@
The weights $w_{1,2}$ and the basis-function coefficients $c_{i}$ are the trainable parameters of the spline. The basis functions $B_{i}$ are chosen as polynomials of degree $k$, with default value $k=3$. The grid parameter $G$ determines how many basis functions build the B-spline and serves as a hyperparameter of the KANs. Furthermore, the domain of the activation function needs to be chosen and can be updated several times during network training to match the input range of the activation function. Specifically, for given parameters $k$, $G$ and the domain $[t_{0},t_{G}]$, a vector $\vec{t}=(t_{-k},\ldots,t_{0},\ldots,t_{G+k})$ of equidistant knot points is constructed. Then, $G+k$ basis functions $B_{i}^{k}(x)$ are recursively defined:
@@ 4 Results
The output layers of all trained models consist of a single node. Although KANs learn activation functions and a learnable function can also be placed on the output node, we choose the sigmoid function to normalize the model outputs to the range $(0,1)$. In addition, for the MLPs, we choose the sigmoid function as output activation. For all trainings, we use the binary cross-entropy as a loss function. We consider $tH$ events as signal (label 1) and $t\bar{t}H$ as background (label 0).
We use the Adam [103] optimizer to train our models. For the KAN trainings, we also compare with the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) [104] optimizer, as it was used in Ref. [17]. For Adam, we find stable trainings for all tested learning rates in the range $[10^{-4},10^{-3}]$ and select $3\times 10^{-4}$ for all trainings. The training is performed in mini-batches of batch size $256$. Trainings with the L-BFGS optimizer are performed using the entire dataset in full-batch training with a learning rate of $10^{-3}$. No significant performance differences between KANs trained with the two optimizers were found and hence we use Adam for all trainings discussed below due to its faster convergence. To ensure that all models converge during training and to allow for a fair performance comparison, we employ early stopping. The loss obtained from the validation dataset is monitored during training and if there is no improvement over 25 epochs, the training is terminated. The model parameters from the epoch with the lowest loss on the validation set are used for the comparisons.
@@
The evolution of the loss and accuracy [^4] The accuracy is defined as the fraction of correctly classified examples with a decision threshold of 0.5 in the network output. over the training is shown in Fig. 3 for three KANs. As expected, the single-layer KAN shows the lowest performance among these models but still reaches an accuracy of $81.5\%$ on the validation set, which is only about $1.5$ percentage points lower than the accuracy of the two-layer and four-layer KANs. The latter two models reach similar accuracies and loss values. The two-layer KAN, with its wide second layer, has the highest number of trainable parameters in this comparison. It converges within the fewest number of training epochs and at the same time shows the largest generalization gap. While the range of the input variables is fixed with the min-max scaling, this is not the case for the input range of subsequent layers in the networks. Therefore, we use the default setting to update the domains ten times within the first 50 epochs of training. Training instabilities are visible at these epochs.
@@
The output distributions of the test dataset classified by these KANs are shown in Fig. 4. The separation of the two classes is clearly visible with the networks accumulating the majority of the events close to the respective label. Also here, only slight differences are visible between the two- and the four-layer network, while the one-layer KAN achieves a visibly worse separation.
A possible advantage of KANs over MLPs lies in their potential for interpretability. While the patterns learned by MLPs are usually embedded in large matrices of trainable parameters and thus hard to understand, multiple trained KAN parameters can be visualized in form of a single spline. As demonstrated in Ref. [17], patterns learned by KANs can often be understood more easily. However, these examples are much lower in dimensionality and complexity than typical HEP machine-learning tasks. In our study with 22 input features, we find that the interpretability of wide models, such as the 22–45–1 KAN, is limited. For instance, this particular model consists of more than $1000$ learnable activation functions. Its visualization is, hence, complex and difficult for humans to interpret. Therefore, we focus on shallow KAN structures for the interpretability.
In Fig. 5, the one-layer KAN is depicted together with its learned activation functions. The corresponding figure for the KAN with a shallow second layer (22–3–1) is shown in the Appendix. In the following, we discuss the interpretability of the learned activation functions.
@@
These values directly indicate the importance of the different input features for the one-layer KAN classification output due to the simple summation of input features transformed by a single activation function on its output node. The feature importances obtained from the commonly used techniques of permutation feature importance and Shapley values yield similar patterns. This confirms that the $L_{1}$-norms in the one-layer KAN effectively capture feature importance. This network is expected to exploit the features with strong discrimination between the two classes. Values of the $L_{1}$-norms close to zero are found for the azimuthal angles, which provide no discrimination power on their own. The largest values of the $L_{1}$-norms are found for the lepton multiplicity, with which the network can identify di-leptonic signatures present in a fraction of the $t\bar{t}H$ events but mostly absent in the $tH$ process, and for variables based on (pseudo-)rapidity differences as well as the jet multiplicity. These reflect the good separation that can be achieved with these variables on their own. In deeper KANs, where multiple learnable functions transform each feature and their outputs are combined in a more complex manner, the $L_{1}$-norms are not directly interpretable as feature importance scores.
@@
In Fig. 6, the learned activation functions are shown for five examples of features. These features were chosen to represent a set of variables with different shapes of their distributions, where the corresponding splines of the single-layer KAN have high $L_{1}$-norms. The distributions of the pre-processed features are also shown for the two classes, together with the univariate log-likelihood ratios of the signal over the background. As expected in binary classification, where the optimal decision boundary is related to the log-likelihood ratio of the classes, the activation functions learned by the one-layer KAN resemble these univariate log-likelihood ratios of the respective input features. This resemblance is anticipated because the one-layer KAN sums univariate functions of each input feature, aligning with the additive form of the multivariate log-likelihood ratio, which decomposes into a sum of univariate log-likelihood ratios when features are independent. However, due to correlations, differences between the learned activation functions and the univariate log-likelihood ratios may arise, as we find, for example, in the normalization of the spline transforming the $\eta(j)$ variable. The corresponding splines obtained from the training of the 22–3–1 KAN are also shown in the same panels. These differ clearly from the likelihood ratios and the 22–1 KAN splines. This is consistent with the expectation that multi-layer KANs learn more abstract representations of the data. As the splines acting on the same input feature are allowed to have different functional shapes or $L_{1}$-norms, each of the three nodes of the internal layer of the 22–3–1 KAN captures different aspects of the data.
@@
Receiver operating characteristic (ROC) curves for selected models are shown in Fig. 8, including models with deep and shallow structures for both, MLPs and KANs. The overall shape of the ROC curves of KANs and MLPs is very similar, except for the one-layer KAN, where the area under the curve (AUC) is considerably lower. We observe that an MLP with only a single hidden layer reaches an AUC [^5] The uncertainty in the quoted AUCs from the limited size of the test dataset are approximately $6\times 10^{-4}$. of $0.906$, only slightly below the AUCs obtained by our best MLP ($0.908$) and the best KAN, where the 22–45–1 network also reaches an AUC of $0.908$. We find only slight differences in classification performance between well-tuned KANs and MLPs.
@@
To evaluate the parameter efficiency, we compare the performance of KANs and MLPs as a function of the number of trainable model parameters. A small number of parameters is favorable for a given performance, as smaller models are computationally more efficient, better interpretable (if at all possible), and less prone to overfitting. In Fig. 8, two metrics often used in HEP for classification performance evaluation are included: the AUC, and the background rejection for a fixed signal efficiency, here chosen as $70\,\%$. The rejection is defined as the inverse of the efficiency for a given threshold on the network output. Overall, we find that for very low numbers of parameters, the MLPs outperform the KANs, while for medium and high numbers of parameters, the performance of KANs and MLPs is similar. The smallest MLP trained with only eight nodes in the hidden layer has only $193$ parameters and has reached a background rejection of $12.9$ already [^6] The uncertainty in the quoted rejection values from the limited size of the test dataset is approximately $0.2$.. A similar value is achieved by the 22–3–1 KAN, which has $556$ trainable parameters. While increasing the number of parameters of the MLPs, their rejection saturates at around $14.9$. The KANs achieve rejections ranging from $11.4$ for the single-layer network to $15.2$ for the 22–45–1 KAN. The deeper KANs from this study show a similar performance. Instead of varying the node count of KANs, the number of model parameters can be increased as well by raising the grid parameter $G$. For illustration, the 22–45–1 KAN is included in Fig. 8 when trained with $G=10$ and $G=25$. For KANs with grid parameters considerably higher than the default of $G=5$, we find that these models overtrain faster and generalize worse.